OpenStudy (anonymous):

The derivative of cos^2(4x) is 2cos(4x)(-sin(4x))(4) Why do I take the derivative of the 4x? Why is that last 4 present?

OpenStudy (raden):

that's how the chain rule works, u have to derive from inner till outer

OpenStudy (raden):

derivative of 4x is 4

OpenStudy (raden):

derivative of (...)^2 = 2(...)

OpenStudy (raden):

derivative of cos() = -sin()

OpenStudy (raden):

times up of them

OpenStudy (anonymous):

Yes, I understand all of the rest of it. But I don't know why I have to take the derivative of the 4x. I guess I just don't understand the chain rule.

OpenStudy (raden):

well, i will show step by step

OpenStudy (raden):

we have y = cos^2(4x) let u = 4x du/dx = 4, agree ?

OpenStudy (anonymous):

yes

OpenStudy (raden):

ok, now we have y = cos^2 (u) let v = cos(u) dv/du = -sin(u), agree ?

OpenStudy (anonymous):

ok

OpenStudy (raden):

finally, now we have y = v^2 dy/dv = 2v, right ?

OpenStudy (anonymous):

right

OpenStudy (raden):

thus, to get y' = dy/dx just multiply of them dy/dx = du/dx * dv/du * dy/dv dy/dx = 4 * (-sin(u)) * 2v

OpenStudy (raden):

but, the answer must be in x term so, subtitute back that u = 4x and v = cos(u) = cos(4x)

OpenStudy (raden):

therefore, dy/dx = 4 * (-sin(u)) * 2v dy/dx = 4 * (-sin(4x)) * 2cos(4x)

OpenStudy (raden):

that's look too long way, so u have to doing the chain rule directly to be easier

OpenStudy (anonymous):

Ok, so cos(x) would actually have a derivative of -sin(x) * 1 where the one is the derivative of x?

OpenStudy (raden):

exactly, yes!

OpenStudy (anonymous):

Thank you very much! You've helped a lot

OpenStudy (raden):

my pleasure could help u :)